OpenStudy (anonymous):

y = 9^(5x + 8) Find the derivative Ans correct? - 9^(5x + 8) ln(9)*(5)

OpenStudy (anonymous):

do you know what the derivative of : \[y = a^x\] is? (a is just a number, like 2, or 3, or whatever)

OpenStudy (anonymous):

x*a-x

OpenStudy (anonymous):

chain rule! :)

OpenStudy (anonymous):

you cant treat it like: and just pull the power down and subtract one, those are different functions. I wont bother you with the proof (unless you want to see it, its not too long), but if: \[y = a^x\] the derivative is: \[y' = a^xln(a)\]

OpenStudy (anonymous):

er, i meant you cant treat it like \[y = x^n\] i wonder where the equation went >.<

OpenStudy (anonymous):

so was the ans correct i wasn't sure how to multiply the last part

OpenStudy (anonymous):

oh, i didnt see you had attempted already, my bad >.< your answer is correct. you can leave it that way, or if you want make the 5 a power on the 9, but it doesnt really matter.

OpenStudy (anonymous):

thanks man